OpenStudy (anonymous):

smallville college has 8 faculty, 6 of whom are tenured. the chairman wants to establish a committee of 3 faculty members to review the curriculum. if she selects the committee at random: a) what is the probability all members of the committee are tenured? b) what is the probability at least one member is not tenured?

pooja195 (pooja195):

a. The total number of ways to choose 3 members out of 8 is C(8, 3) = 8!/(3! 5!) = (8*7*6*5*4*3*2*1) / ((3*2*1) * (5*4*3*2*1)) =? The total number of ways to choose 3 tenured people from 6 is C(6,3) = 6!/(3! 3!) = (6*5*4*3*2*1) / ((3*2*1) * (3*2*1))=? This is not my work i found it here: http://www.justanswer.com/calculus-and-above/6i4nu-computer-systems-department-eight-faculty-six-whom.html I wanted to help you but I did not know this stuff so once again this is NOT my work.